graph the linear equation. find threee points that solve the equation, then plot on the graph. -x - 2y= -10
Found 2 solutions by Fombitz, ewatrrr:
Answer by Fombitz(32388)

When x=0,-2y=-10,y=5
When y=0, -x=-10,x=10
When x=2, -2-2y=-10,-2y=-8,y=4
.
.
.

.
.
.
Then draw the line through the points.


Answer by ewatrrr(24785)

 
Hi
-x - 2y= -10
y = -x/2 + 5
x y
0 5
2 4
4 3
plot and connect with line